find two consecqutive intigers whos product is 182.
Thank you

Answer by stanbon(48579)

find two consecutive integers whose product is 182.
x(x+1)=182
x^2+x-182=0
(x+14)(x-13)=0
x=-14 or x=13
Solutions are -14,-13 OR 13,14
Cheers,
Stan H.
Cheers,